OpenStudy (kaliyah0524):

Will give medal! solve 3(x + 1) + 6 = 33 I got 8 but the choices are a. 15 b. 15/2 c. 9 d. 11

OpenStudy (bibby):

2 different ways of going about this. 1) distributing the 3, and solving for x 2) subtracting 6 from both sides and then dividing both sides by 3

OpenStudy (kaliyah0524):

nevermind I wrote the wrong problem . it's x: 1 over 4 (2x − 14) = 4 that I got 8 for

OpenStudy (calculusxy):

3(x + 1) + 6 = 33 3x + 3 + 6 = 33 3x + 9 = 33 <--- I combined like terms (3 and 6) 3x + 9 - 9 = 33 - 9 3x = 24 3x/3 = 24/3 <---- you want to isolate x so you divide both sides by 3 x = 8

OpenStudy (bibby):

\(\cancel{4}\dfrac{1}{\cancel{4}}(2x-14)=4(4)\\2x-14=16\\2x=30\\x=15\)

OpenStudy (kaliyah0524):

Thankyou!